# A Multi-Component, Multi-Physics Computational Model for Solving Coupled Cardiac Electromechanics and Vascular Haemodynamics

### v1

@@ Acknowledgements
We acknowledge funding support from the European Commission CompBioMed Centre of Excellence (Grant No. 675451 and 823712) and the UK Engineering and Physical Sciences Research Council (EPSRC) under the projects “UK Consortium on Mesoscale Engineering Sciences (UKCOMES)” (Grant No. EP/R029598/1), “Software Environment for Actionable & VVUQ-evaluated Exascale Applications (SEAVEA)” (Grant No. EP/W007711/1), and “CompBioMedX: Computational Biomedicine at the Exascale” (Grant No. EP/X019446/1). S. C. Y. L. is grateful for the research studentship funded by University College London and CBK Sci Con Ltd. A. Z. and M. V. acknowledge the European Union — EIC Project No 190134524: “ELEM Virtual Heart Populations for Supercomputers” (ELVIS). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or EISMEA. Neither the European Union nor the granting authority can be held responsible for them. This work used the ARCHER2 UK National Supercomputing Service (https://www.archer2.ac.uk), made accessible via the EPSRC CompBioMedX project, and resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.

### v2

@@ Acknowledgements
We acknowledge funding support from the European Commission CompBioMed Centre of Excellence (Grant No. 675451 and 823712) and the UK Engineering and Physical Sciences Research Council (EPSRC) under the projects “UK Consortium on Mesoscale Engineering Sciences (UKCOMES)” (Grant No. EP/R029598/1), “Software Environment for Actionable & VVUQ-evaluated Exascale Applications (SEAVEA)” (Grant No. EP/W007711/1), and “CompBioMedX: Computational Biomedicine at the Exascale” (Grant No. EP/X019446/1). S. C. Y. L. is grateful for the research studentship funded by University College London and CBK Sci Con Ltd. A. Z. and M. V. acknowledge support from the European Union through the EIC Project No. 190134524, “ELEM Virtual Heart Populations for Supercomputers” (ELVIS), as well as the Horizon Europe research and innovation programme under grant agreement No. 101136728 (VITAL). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or EISMEA. Neither the European Union nor the granting authority can be held responsible for them. This work used the ARCHER2 UK National Supercomputing Service (https://www.archer2.ac.uk), made accessible via the EPSRC SEAVEA project, and resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725.